OpenStudy (anonymous):

The length of the two legs of a triangle are 13 centimeters and 10 centimeters .what is the exact length of the hypotenuse ?

OpenStudy (anonymous):

Pythagorean theorem says a^2 + b^2 = c^2, c being the hypotenuse. We know a and b. Plug those in. 13^2 + 10^2 = c^2. 169 + 100 = c^2. c^2 = 269. c = sqrt 269. c = 16.40.